# Stochastic Sequential Quadratic Programming for Optimization with Functional Constraints

### v1

@@ Abstract
Stochastic convex optimization problems with nonlinear functional constraints are ubiquitous in machine learning applications, including multi-task learning, structured prediction, and multi-view learning. The presence of non-linear functional constraints renders the traditional projected stochastic gradient descent and related projection-based methods inefficient, and motivates the use of first-order methods. However, existing first-order methods, including primal and primal–dual algorithms, typically rely on a bounded (sub-)gradient assumption, which may be too restrictive in many settings.
We propose a stochastic sequential quadratic programming (SSQP) algorithm that works entirely in the primal domain, avoids projecting onto the feasible region, obviates the need for bounded gradients, and achieves state-of-the-art oracle complexity under standard smoothness and convexity assumptions. A faster version, namely SSQP-Skip, is also proposed where the quadratic subproblems can be skipped in most iterations. Finally, we develop an accelerated variance-reduced version of SSQP (VARAS), whose oracle complexity bounds match those for solving unconstrained finite-sum convex optimization problems. The superior performance of the proposed algorithms is demonstrated via numerical experiments on real datasets.
@@ I Introduction
where $f({\mathbf{x}}):=\mathbb{E}_{t}[f_{i_{t}}({\mathbf{x}})]$ and $\mathbb{E}_{t}[\cdot]$ denotes the expectation with respect to the random index $i_{t}$. We will also consider a finite-sum case, which arises when the index $i_{t}$ is sampled uniformly from $\{1,\ldots,n\}$, so that $f({\mathbf{x}}):=\frac{1}{n}\sum_{i=1}^{n}f_{i}({\mathbf{x}})$. The functions $f_{i}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ and $g_{k}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ are proper, closed, convex, and $L$-smooth. The regularization function $h:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ is convex but possibly non-smooth, and may include an indicator function corresponding to a set-inclusive constraint of the form ${\mathbf{x}}\in{\mathcal{K}}$ for a closed convex set $\mathcal{K}$}. The stochastic objective function in ($\mathcal{P}$) commonly arises in the context of stochastic approximation and online learning. The finite-sum structure in particular is widely used in supervised learning tasks where the goal is to minimize the empirical risk over $n$ training samples. Non-linear functional constraints similarly arise in many learning tasks, such as multi-task learning, where the constraints capture the relationships between tasks [1], structured prediction, where the constraints may encode structural prior knowledge about the feature space [2], and semi-supervised multi-view object recognition, where the constraints model the relationship between views [3].
We consider the high-dimensional setting, where for a given $\mathbf{x}$}, a stochastic first-order oracle (SFO) provides us with $\nabla f_{i}({\mathbf{x}})$ for a randomly selected index $i$ as well as $\{g_{k}({\mathbf{x}}),\nabla g_{k}({\mathbf{x}})\}_{k=1}^{m}$. In such settings, classical methods such as the projected stochastic gradient descent (SGD) lose their efficacy, since each iteration requires projection onto a feasible region defined by functional constraints, which itself is a complicated operation. Instead, efficient and scalable algorithms for solving ($\mathcal{P}$) must rely only on the first-order information provided by the oracle. The SFO complexity of an algorithm is defined as the number of SFO calls required to achieve an $\epsilon$}-optimal solution, which may be a random vector ${\mathbf{x}}\in{\mathbb{R}}^{d}$ such that
@@
This work puts forth a new class of stochastic sequential quadratic programming (SSQP) algorithms for solving ($\mathcal{P}$). The key idea behind the proposed approach is to reformulate ($\mathcal{P}$) as an unconstrained non-smooth optimization problem using the exact penalty method [16, 17] and to solve it via the prox-linear algorithm [18] using only stochastic first-order information. In the present case, each iteration reduces to solving a diagonal quadratic program (QP), instead of a full projection or general convex subproblems, while still achieving (near-)optimal SFO complexity and avoiding any bounded (sub-)gradient assumptions. Our main algorithmic contributions are: (i) a vanilla SSQP algorithm with SFO complexity matching state-of-the-art primal–dual algorithms for solving ($\mathcal{P}$); (ii) an SSQP-Skip algorithm that solves the proximal and QP subproblems only infrequently, while retaining the same SFO complexity guarantees; and (iii) a first accelerated variance-reduced algorithm, VARAS, for finite-sum constrained optimization problems, whose SFO complexity is nearly on par with the best known methods in the unconstrained setting. Central to these algorithms is a novel one-step inequality that leads to the required optimality gap and constraint violation bounds. The numerical performance of the proposed variants is also tested on two real-world problems and substantially outperforms representative baselines for solving ($\mathcal{P}$). The proposed algorithms differ fundamentally from classical sequential QP methods, which require full gradient and Hessian information at each iteration [19, Sec. 4.3.1] and are therefore ill-suited to large-scale stochastic settings; in contrast, the proposed SSQP methods operate with only stochastic first-order information.
@@ I-A Related Work
The exact penalty reformulation of constrained problems has been well studied in convex optimization, and a sequential QP approach is detailed in [19, Sec. 4.3.1]. The objective of the exact-penalty reformulation can be viewed as a compositional optimization problem, and in the deterministic case it can be solved using the approach of [22]. The corresponding stochastic problems can be addressed using the model-based framework of [23] or the prox-linear method of [18]. A different exact-penalty method was proposed in [24], where the focus is on non-convex equality constraints, leading to an oracle complexity of $\mathcal{O}({\epsilon}^{-3.5})$ for finding an $\epsilon$}-stationary point. Using smooth approximations of exact penalties, so as to work with their gradients, is another alternative but typically yields weaker rates [25, 26]. Our work can be viewed as bringing these exact-penalty and prox-linear ideas into a stochastic SQP framework with explicit oracle-complexity guarantees.
@@
We remark that among these, [9, 4, 39, 6] adopt a different SFO model for ($\mathcal{P}$), wherein at iteration $t$, only $\{g_{j_{t}}({\mathbf{x}}),\nabla g_{j_{t}}({\mathbf{x}})\}$ for random index $j_{t}$, is revealed. While this greatly reduces the per-iteration cost, these algorithms require strong regularity assumptions that couple individual constraint violations with the distance to the full feasible set (see, e.g., [4, Assumption 4]), as well as bounded subgradients of both the objective and the constraints. Their resulting convergence rates match those of stochastic subgradient methods and depend explicitly on the regularity constant of the constraint system. In contrast, the proposed SSQP method uses all functional constraint gradients at every iteration, and therefore does not rely on such regularity conditions, while attaining (near-)optimal oracle complexity. Nevertheless, a single-constraint variant of SSQP with improved rates may be an interesting direction for future work.
@@
SQP ${\mathcal{O}}\left({\sqrt{\frac{n}{\epsilon}}}\right)$ ${\mathcal{O}}\left({n\log n+\sqrt{n}\log\frac{1}{{\epsilon}}}\right)$
@@ II-B Exact Penalty Reformulation
We reformulate the problem using the exact penalty method [19, Sec. 4.3.1], so as to obtain:
@@ III-A SSQP Algorithm
for $t\geq 0$, where $i_{t}$ is a random index. Observe that when ${\mathcal{K}}={\mathbb{R}}^{d}$, the updates bear resemblance to the sequential quadratic programming approach proposed in [47, Sec. 4.3.1], and hence we refer to our algorithm, summarized in Algorithm 1, as Stochastic SQP (SSQP).
@@ III-B SSQP-Skip algorithm
The result in Theorem 2 is the first of its kind in the context of constrained optimization, and the proof is provided in Appendix C. The bound in Theorem 2 is even better than that obtained for the corresponding unconstrained problem in [50]. Specifically, the proof of Algorithm 2 proceeds in a similar manner and obtains a similar recursion as that in [50, Lemma C.2]. However, we utilize diminishing stepsizes to avoid the $\log(T)$ term that appears in [50, Corollary 5.6].
@@ IV Variance-Reduced Accelerated SSQP Algorithm
The proof of Theorem 3 begins by deriving a one-step inequality based on the update in (26) and the relationships among the parameters. The resulting inequality (see the supplementary material) matches the form of [31, Lemma 6], so the remaining arguments in [31, Theorems 1–2] apply directly. Though the rates established in Theorem 3 are the fastest, the most common case is when $\epsilon$ is small and VARAS achieves SFO complexity of $\mathcal{O}\left(\frac{1}{\sqrt{\epsilon}}\right)$} and $\mathcal{O}\left(\log\left({\frac{1}{\epsilon}}\right)\right)$} for the convex and strongly convex cases, respectively. These rates are clearly superior to the best known rates for constrained problems in [54, 55, 56]. Before concluding the theoretical results, the following remark is due.
@@ Appendix C Proof of Theorem 2
Before establishing the main result, we derive some intermediate results through the following lemmas. The following result is a consequence of the update in (19) and the convexity and smoothness of $g_{k}$.
@@
Proof:
[…]
$+\gamma\max\left\{{\left[{g_{k}({\tilde{{\mathbf{x}}}}_{t+1})+\langle\nabla g_{k}({\tilde{{\mathbf{x}}}}_{t+1}),{\mathbf{x}}_{\star}-{\tilde{{\mathbf{x}}}}_{t+1}\rangle}\right]_{+}}\right\}.$ (73)
Finally, the first order optimality condition of ($\mathcal{P}_{1}$) implies that
where (77) follows from the convexity of $g_{k}$ and monotonicity of the $\max\{[\cdot]_{+}\}$ operator. Substituting (77) in (C) and multiplying by $2\eta_{t}$, we obtain the required result. ∎
@@
We also note from the update of ${\mathbf{y}}_{t+1}$ in Algorithm 2 that
${\mathbf{y}}_{t+1}={\mathbf{y}}_{t}+\tfrac{w_{t}p_{t}}{2\eta_{t}}({\hat{{\mathbf{x}}}}_{t+1}-{\tilde{{\mathbf{x}}}}_{t+1})$ (80)
so that
@@
Substituting (C)-(85) in (83), we obtain the required result. ∎
@@
The key to proving the required result is the following one-step inequality, which looks similar to the result in [31, Lemma 6] but requires a different proof from that in the proximal case.
@@ C-A Proof of Theorem 3(1)
so that ${\mathcal{L}}_{s}{\tilde{\Delta}}_{s}={\mathcal{R}}_{s}{\tilde{\Delta}}_{s-1}+\tfrac{1}{2}(\delta^{z}_{T_{s-1}}-\delta^{z}_{T_{s}})$. Summing over $j=1,..,s$ and rewriting,
${\mathcal{L}}_{s}\Delta_{s}+\sum_{j=1}^{s-1}\left({{\mathcal{L}}_{j}-{\mathcal{R}}_{j+1}}\right){\tilde{\Delta}}_{j}\leq{\mathcal{R}}_{1}{\tilde{\Delta}}_{0}+\tfrac{1}{2}\delta^{z}_{0}-\tfrac{1}{2}\delta^{z}_{T_{s}}$ (122)
where the last term can be dropped.
@@ C-B2 Case $s>s_{0}$ and $n\geq\frac{3\kappa}{4}$
$N_{\text{SFO}}$ $=nS+\sum_{s=1}^{S}T_{s}\leq 2nS={\mathcal{O}}\left({n\log\tfrac{D_{0}}{{\epsilon}}}\right).$ (141)
$N_{\text{QMO}}$ $=\sum_{s=1}^{S}T_{s}={\mathcal{O}}\left({n+(S-s_{0})n}\right)={\mathcal{O}}\left({n\log\tfrac{D_{0}}{{\epsilon}}}\right)$ (142)
@@ C-B4 Case $s>s_{0}+\sqrt{\frac{12\kappa}{n}}-4$ and $n<\frac{3\kappa}{4}$
As in [31, Lemma 11], we can establish that $\sum_{t=1}^{T_{s_{0}}}\theta_{t}\geq\Gamma_{T_{s_{0}}}\Omega_{s_{0}}$ where

### v2

@@ Abstract
Stochastic convex optimization problems with nonlinear functional constraints are ubiquitous in signal processing applications including constrained least-squares, set-membership adaptive filtering, and trajectory optimization under uncertain fields. The presence of non-linear functional constraints renders the traditional projected stochastic gradient descent and related projection-based methods inefficient, and motivates the use of first-order methods. However, existing first-order methods, including primal and primal–dual algorithms, typically rely on a bounded (sub-)gradient assumption, which may be too restrictive in high-dimensional settings. We propose a stochastic sequential quadratic programming (SSQP) algorithm that works entirely in the primal domain, avoids projecting onto the feasible region, obviates the need for bounded gradients, and achieves state-of-the-art oracle complexity under standard smoothness and convexity assumptions. A faster version, namely SSQP-Skip, is also proposed where the quadratic subproblems can be skipped in most iterations. Finally, we develop an accelerated variance-reduced version of SSQP (VARAS), whose oracle complexity bounds match those for solving unconstrained finite-sum convex optimization problems. The superior performance of the proposed algorithms is demonstrated via numerical experiments on real datasets.
@@ I Introduction
where $f({\mathbf{x}}):=\mathbb{E}_{t}[f_{i_{t}}({\mathbf{x}})]$ and $\mathbb{E}_{t}[\cdot]$ denotes the expectation with respect to the random index $i_{t}$. We will also consider a finite-sum case, which arises when the index $i_{t}$ is sampled uniformly from $\{1,\ldots,n\}$, so that $f({\mathbf{x}}):=\frac{1}{n}\sum_{i=1}^{n}f_{i}({\mathbf{x}})$. The functions $f_{i}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ and $g_{k}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ are proper, closed, convex, and $L$-smooth. The regularization function $h:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ is convex but possibly non-smooth, and may include an indicator function corresponding to a set-inclusive constraint of the form ${\mathbf{x}}\in{\mathcal{K}}$ for a closed convex set $\mathcal{K}$}. The stochastic objective function in ($\mathcal{P}$) commonly arises in the context of stochastic approximation and online signal processing tasks such as streaming estimation [1]. The finite-sum structure in particular is widely used in batch estimation, system identification, and empirical risk formulations built from a dataset with $n$ training samples [2]. Non-linear functional constraints similarly arise in many signal processing tasks. For example, in trajectory generation and navigation under uncertain environmental fields [3], constraints encode vehicle kinematics and actuator limits. Similarly, in set membership adaptive filtering, constraints can enforce residual bounds for critical samples [4, 5, 6].
We consider the high-dimensional setting, where for a given $\mathbf{x}$}, a stochastic first-order oracle (SFO) returns a stochastic gradient $\nabla f_{i_{t}}({\mathbf{x}})$, e.g., from a randomly selected data sample or snapshot, together with constraint values and gradients $\{g_{k}({\mathbf{x}}),\nabla g_{k}({\mathbf{x}})\}_{k=1}^{m}$. In such settings, classical methods such as the projected stochastic gradient descent (SGD) lose their efficacy, since each iteration requires projection onto a feasible region defined by functional constraints, an operation that is computationally expensive in high-dimensional signal processing models, and often impractical when constraints encode dynamics or per-sample performance requirements. Instead, efficient and scalable algorithms for solving ($\mathcal{P}$) must rely only on the first-order information provided by the oracle. The SFO complexity of an algorithm is defined as the number of SFO calls required to achieve an $\epsilon$}-optimal solution, which may be a random vector ${\mathbf{x}}\in{\mathbb{R}}^{d}$ such that
@@
This work puts forth a new class of stochastic sequential quadratic programming (SSQP) algorithms for solving ($\mathcal{P}$). The key idea behind the proposed approach is to reformulate ($\mathcal{P}$) as an unconstrained non-smooth optimization problem using the exact penalty method [19, 20] and to solve it via the prox-linear algorithm [21] using only stochastic first-order information. In the present case, each iteration reduces to solving a diagonal quadratic program (QP), instead of a full projection or general convex subproblems, while still achieving (near-)optimal SFO complexity and avoiding any bounded (sub-)gradient assumptions. Our main algorithmic contributions are: (i) a vanilla SSQP algorithm with SFO complexity matching state-of-the-art primal–dual algorithms for solving ($\mathcal{P}$); (ii) an SSQP-Skip algorithm that solves the proximal and QP subproblems only infrequently, while retaining the same SFO complexity guarantees; and (iii) a first accelerated variance-reduced algorithm, VARAS, for finite-sum constrained optimization problems, whose SFO complexity is nearly on par with the best known methods in the unconstrained setting. Central to these algorithms is a novel one-step inequality that leads to the required optimality gap and constraint violation bounds. The numerical performance of the proposed variants is also tested on two standard signal processing problems and substantially outperforms representative baselines for solving ($\mathcal{P}$). The proposed algorithms differ fundamentally from classical sequential QP methods, which require full gradient and Hessian information at each iteration [22, Sec. 4.3.1] and are therefore ill-suited to large-scale stochastic settings; in contrast, the proposed SSQP methods operate with only stochastic first-order information. Beyond the stochastic or finite-sum settings, the proposed framework is also compatible with real-time or streaming regimes, where its tracking performance can be studied using similar tools [23].
@@ I-A Related Work
The exact penalty reformulation of constrained problems has been well studied in convex optimization, and a sequential QP approach is detailed in [22, Sec. 4.3.1]. The objective of the exact-penalty reformulation can be viewed as a compositional optimization problem, and in the deterministic case it can be solved using the approach of [26]. The corresponding stochastic problems can be addressed using the model-based framework of [27] or the prox-linear method of [21]. A different exact-penalty method was proposed in [28], where the focus is on non-convex equality constraints, leading to an oracle complexity of $\mathcal{O}({\epsilon}^{-3.5})$ for finding an $\epsilon$}-stationary point. Using smooth approximations of exact penalties, so as to work with their gradients, is another alternative but typically yields weaker rates [29, 30]. Our work can be viewed as bringing these exact-penalty and prox-linear ideas into a stochastic SQP framework with explicit oracle-complexity guarantees.
@@
We remark that among these, [12, 7, 43, 9] adopt a different SFO model for ($\mathcal{P}$), wherein at iteration $t$, only $\{g_{j_{t}}({\mathbf{x}}),\nabla g_{j_{t}}({\mathbf{x}})\}$ for random index $j_{t}$, is revealed. While this greatly reduces the per-iteration cost, these algorithms require strong regularity assumptions that couple individual constraint violations with the distance to the full feasible set (see, e.g., [7, Assumption 4]), as well as bounded subgradients of both the objective and the constraints. Their resulting convergence rates match those of stochastic subgradient methods and depend explicitly on the regularity constant of the constraint system. In contrast, the proposed SSQP method uses all functional constraint gradients at every iteration, and therefore does not rely on such regularity conditions, while attaining (near-)optimal oracle complexity. Nevertheless, a single-constraint variant of SSQP with improved rates may be an interesting direction for future work.
@@
SQP ${\mathcal{O}}\left({\sqrt{\frac{n}{\epsilon}}}\right)$ ${\mathcal{O}}\left({n\log n+\sqrt{n}\log\tfrac{1}{{\epsilon}}}\right)$
@@ II-B Exact Penalty Reformulation
We reformulate the problem using the exact penalty method [22, Sec. 4.3.1], so as to obtain:
@@ III-A SSQP Algorithm
for $t\geq 0$, where $i_{t}$ is a random index. Observe that when ${\mathcal{K}}={\mathbb{R}}^{d}$, the updates bear resemblance to the sequential quadratic programming approach proposed in [51, Sec. 4.3.1], and hence we refer to our algorithm, summarized in Algorithm 1, as Stochastic SQP (SSQP).
@@ III-B SSQP-Skip algorithm
The result in Theorem 2 is the first of its kind in the context of constrained optimization, and the proof is provided in Appendix C. The bound in Theorem 2 is even better than that obtained for the corresponding unconstrained problem in [54]. Specifically, the proof of Algorithm 2 proceeds in a similar manner and obtains a similar recursion as that in [54, Lemma C.2]. However, we utilize diminishing stepsizes to avoid the $\log(T)$ term that appears in [54, Corollary 5.6].
@@ IV Variance-Reduced Accelerated SSQP Algorithm
The proof of Theorem 3 begins by deriving a one-step inequality based on the update in (26) and the relationships among the parameters. The resulting inequality (see the supplementary material) matches the form of [35, Lemma 6], so the remaining arguments in [35, Theorems 1–2] apply directly. Though the rates established in Theorem 3 are the fastest, the most common case is when $\epsilon$ is small and VARAS achieves SFO complexity of $\mathcal{O}\left(\frac{1}{\sqrt{\epsilon}}\right)$} and $\mathcal{O}\left(\log\left({\frac{1}{\epsilon}}\right)\right)$} for the convex and strongly convex cases, respectively. These rates are clearly superior to the best known rates for constrained problems in [58, 59, 60]. Before concluding the theoretical results, the following remark is due.
@@ Appendix C Proof of Theorem 2
Before establishing the main result, we derive some intermediate results through the following lemmas. The following result is a consequence of the update in (19) and the convexity and smoothness of $g_{k}$ and its proof is provided in the Supplementary material.
@@
We also note from the update of ${\mathbf{y}}_{t+1}$ in Algorithm 2 that ${\mathbf{y}}_{t+1}={\mathbf{y}}_{t}+\tfrac{w_{t}p_{t}}{2\eta_{t}}({\hat{{\mathbf{x}}}}_{t+1}-{\tilde{{\mathbf{x}}}}_{t+1})$ so that
@@
Substituting (C)-(79) in ($7$), we obtain the required result. ∎
@@
Proof of Lemma 3
Proof:
[…]
$\hskip-2.84526pt+\gamma\max\left\{{\left[{g_{k}({\tilde{{\mathbf{x}}}}_{t+1})+\langle\nabla g_{k}({\tilde{{\mathbf{x}}}}_{t+1}),{\mathbf{x}}_{\star}-{\tilde{{\mathbf{x}}}}_{t+1}\rangle}\right]_{+}}\right\}.$ (89)
Finally, the first order optimality condition of ($\mathcal{P}_{1}$) yields
where (93) follows from the convexity of $g_{k}$ and monotonicity of the $\max\{[\cdot]_{+}\}$ operator. Substituting (93) in (Proof of Lemma 3) and multiplying by $2\eta_{t}$, we obtain the required result. ∎
@@
The key to proving the required result is the following one-step inequality, which looks similar to the result in [35, Lemma 6] but requires a different proof from that in the proximal case.
@@ C-A Proof of Theorem 3(1)
so that ${\mathcal{L}}_{s}{\tilde{\Delta}}_{s}={\mathcal{R}}_{s}{\tilde{\Delta}}_{s-1}+\tfrac{1}{2}(\delta^{z}_{T_{s-1}}-\delta^{z}_{T_{s}})$. Summing over $j=1,..,s$ and rewriting, we obtain ${\mathcal{L}}_{s}\Delta_{s}+\sum_{j=1}^{s-1}\left({{\mathcal{L}}_{j}-{\mathcal{R}}_{j+1}}\right){\tilde{\Delta}}_{j}\leq{\mathcal{R}}_{1}{\tilde{\Delta}}_{0}+\tfrac{1}{2}\delta^{z}_{0}-\tfrac{1}{2}\delta^{z}_{T_{s}}$ where the last term can be dropped.
@@ C-B2 Case $s>s_{0}$ and $n\geq\frac{3\kappa}{4}$
$N_{\text{SFO}}$ $=nS+\sum_{s=1}^{S}T_{s}\leq 2nS={\mathcal{O}}\left({n\log\tfrac{D_{0}}{{\epsilon}}}\right).$
$N_{\text{QMO}}$ $=\sum_{s=1}^{S}T_{s}={\mathcal{O}}\left({n+(S-s_{0})n}\right)={\mathcal{O}}\left({n\log\tfrac{D_{0}}{{\epsilon}}}\right)$
@@ C-B4 Case $s>s_{0}+\sqrt{\frac{12\kappa}{n}}-4$ and $n<\frac{3\kappa}{4}$
As in [35, Lemma 11], we can establish that $\sum_{t=1}^{T_{s_{0}}}\theta_{t}\geq\Gamma_{T_{s_{0}}}\Omega_{s_{0}}$ where